OpenStudy (anonymous):

Find the product. (x - 7)2 x2 - 7x + 49 x2 - 14x + 49 x2 + 49 x2 - 49

OpenStudy (pfenn1):

\[(x-7)^2=(x-7)(x-7)=x^2-14x+49\]